Hospital records show 12% of all patients are admitted for surgical treatments, 16% are
admitted for obstetrics, and 2% receive both obstetrics and surgical treatment. If a new patient
is admitted to the hospital, what is the probability that the patient will be admitted either for
surgery, obstetrics, or both?
1
Expert's answer
2020-08-09T17:54:33-0400

Solution


Let X represent the reason a patient is admitted


Let 'S' be Patients admitted for surgical treatment, 'O' for obstetrics and 'B' for both

"X=\{S, O, B\}"


"S= 12\%"

"O = 16\%"

"B=2\%"


Probability of being admitted for the different reasons:


"P(X=S) = 0.12"

"P(X=O) = 0.16"

"P(X=B) = 0.02"


P(X=S) + P(X=O) - P(X=B)

"= 0.12 + 0.16 - 0.02 = 0.26"


answer: 0.26